what is the height of a cyliner that has a volume of 5024 and the diameter of 16

Answer by mananth(16946)

what is the height of a cyliner that has a volume of 5024 and the diameter of 16

volume of cylinder = pi*r^2*h
5024=pi*16^2*h
5024/pi*256 = h
h= 6.246
Ananth